OpenStudy (anonymous):

I NEED HELP & ANSWER STRAIGHT UP & EXPLANATION THATS EASY TO UNDERSTAND THANK YOU.

OpenStudy (anonymous):

I NEED HELP LIKE BIG TIME ONLY QUESTION THATS KEEPING ME AWAY FROM A 100% :(

OpenStudy (anonymous):

ok first of all, it is clear or not that \[\sqrt{a^2b^4}=ab^2\]?

OpenStudy (anonymous):

YES B/C YOU BREAK THEM APART & THATS WHAT ITS LEFT

OpenStudy (anonymous):

good okay, now it is also clear that if \(\sqrt{a}=c\sqrt{d}\) that if you square you get the equality \[a=c^2d\]?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

ok good, because you see you have no \(a\) in your possible answers, so we know we can replace \(a\) by \(c^2d\)

OpenStudy (anonymous):

at first the answer that i though it was was C but i dont know yet though but i get you

OpenStudy (anonymous):

also \[\sqrt{b}=d\sqrt{c}\] squaring give \[b=d^2c\] and so also \[b^2=d^4c^2\] ok so far? almost done

OpenStudy (anonymous):

start with \[ab^2\] replace \(a\) by \(c^2d\) and replace \(b^2\) by \(d^4c^2\)

OpenStudy (anonymous):

you get \[\large ab^2=c^2d\times d^4c^2=c^4d^5\]

OpenStudy (anonymous):

wait yes i get you now! & can you like say that